Suppose a​ ten-year, $ 1000 bond with an 8.2 % coupon rate and semiannual coupons is trading for $ 1 035.55.

a. What is the​ bond's yield to maturity​ (expressed as an APR with semiannual​ compounding)?

b. If the​ bond's yield to maturity changes to 9.6 % ​APR, what will be the​ bond's price?

(please show all work)!
Answer #1

a)

Coupon = (8.2% of 1000) / 2 = 41

Number of periods = 10 * 2 = 20

Bond's yield to maturity = 7.68%

Keys to use in a fianncial calculator:

2nd P/Y 2

FV 1000

PV - 1035.55

PMT 41

N 20

CPT I/Y

b)

Rate = 9.6% / 2 = 4.8%

Bond's price = Coupon * [1 - 1 / (1 +r)n] / r + FV / (1 + r)n

Bond's price = 41 * [1 - 1 / (1 +0.048)20] / 0.048 + 1000 / (1 + 0.048)20

Bond's price = 41 * [1 - 0.391538] / 0.048 + 391.538387

Bond's price = 41 * 12.676292 + 391.538387

Bond's price = $911.27